Claim: Fewer than

94​%

of adults have a cell phone. In a reputable poll of

1086

​adults,

85​%

said that they have a cell phone. Find the value of the test statistic.

The value of the test statistic is

nothing.
Answer #1

Test statistic formula for this test is given as below:

Z = (p̂ - p)/sqrt(pq/n)

Where, p̂ = Sample proportion, p is population proportion, q = 1 - p, and n is sample size

n = sample size = 1086

p̂ = x/n = 0.85

p = 0.94

q = 1 - p = 0.06

Z = (p̂ - p)/sqrt(pq/n)

Z = (0.85 - 0.94)/sqrt(0.94*0.06/1086)

Z = -12.4887

Test statistic = -12.4887